John was driving down the highway at 70 mph when he spotted a highway patrolman. Not being a rich kid, he decided to brake to 55 mph. What was the per centage change in his speed.
Answer by RAY100(1637)

(70-55)/70 = 15/70 = .2143 = 21.43 %